# Isocurvature Induced Gravitational Waves at Pulsar Timing Arrays

### v1

@@ 3 Isocurvature Induced Gravitational Waves
$f(\tau,q,|\mathbf{k}-\mathbf{q}|)$ $=$ $\frac{1}{2}\Phi(q,\tau)\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{1}{2}\Psi(q,\tau)\Psi(|\mathbf{k}-\mathbf{q}|,\tau)$ (3.5)
$+$ $\frac{1}{2}\left(\Phi(q,\tau)+\frac{\Psi^{\prime}(q,\tau)}{\cal H}\right)\left(\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{\Psi^{\prime}(|\mathbf{k}-\mathbf{q}|,\tau)}{\cal H}\right)~.$
Here $c$ and $\Psi,\Phi$ are defined in Eq. (2.9) and initial conditions for $\Psi,\Phi$ are given in section 2.
The energy density of gravitational waves relevant for observations is given by [Domenech:SIGWRe]
@@
where $f(x,u,v)$ is given in Eq. (3.5). Plugging in the Green’s function of gravitational waves, the kernel function can be further split as
@@
where $h\equiv H_{0}/(100\,\textrm{km/s/Mpc})$, $g_{*}(T)$ and $g_{*S}(T)$ are the effective number of relativistic degrees of freedom for energy and entropy, respectively. $T_{0}$ is the present-day temperature and $T_{\rm GW}$ is the temperature when the GWs are generated, which is $T_{\rm GW}=O(100)$MeV in this study. Therefore, $g_{*}(T_{\rm GW})=g_{*S}(T_{\rm GW})=17.25$, which includes the contribution from pions, muons, and electrons/positions. The radiation density today $\Omega_{r,0}h^{2}=4.2\times 10^{-5}$ is taken from [Planck:2018vyg].
To explicitly view the effects on induced GWs from different initial conditions, we should look at the source term, in particular $f(x,u,v)$ defined in Eq. (3.5) which encodes all the time evolution. As can be seen from figure 2, the subhorizon dynamics are rather universal, which exhibit damped oscillations. However, the superhorizon behavior depends on the certain initial condition. Both the adiabatic and DRDI cases remain constant at superhorizon $x\ll 1$, whereas the CDI case exhibits a $\tau^{2}$ dependence. These features originate from the fact that $\Phi$ and $\Psi$ are constant for AD and DRDI but $\propto\tau$ for CDI outside the horizon. We note that the CDI case is further suppressed by $(k_{\rm eq}/k)^{2}$ and the DRDI case by $R_{\rm dr}^{2}$, which reflects the suppression of the density of the corresponding species. Therefore, for modes relevant for PTA observations ($k\gg k_{\rm eq}$), the source term of CDI is highly suppressed but that of DRDI can still be sizable for appropriate $R_{\rm dr}^{2}$. The cases of BDI and NDI are similar to CDI with an overall rescaling, and thus we don’t show them here.
@@
Here we would like to clarify some intuitions for curvature and isocurvature induced GWs. One common intuition of SIGWs is that GWs are always sourced by curvature perturbations. In the case of isocurvature initial conditions, the curvature perturbation $\zeta$ starts at zero (in the limit $\tau\to 0$) by definition, but grows and peaks around the horizon crossing. Therefore isocurvature induced GWs are mostly generated around the horizon crossing where the curvature perturbation peaks and the density of induced GWs is also determined by the size of curvature perturbations: $\Omega_{\rm GW}\propto\zeta^{4}|_{k\tau\sim 1}$.
We argue that this intuition is only parametrically correct, but may obscure some interesting features in certain cases. We emphasize that it is the source term in Eq. (3.4) that generates GWs, which is a non-trivial function involving the combination of $\Psi$ and $\Phi$, and their derivatives. Therefore, it is not just curvature perturbations as the source. For the cases without anisotropic stress (including AD, BDI, CDI and NDI before neutrino decoupling), $\Phi=\Psi$ and they have a similar time dependence as $\zeta$ (see figure 1). Therefore, the source term can be roughly rewritten in terms of the curvature perturbation $\zeta$. This explains the suppression of $\Omega_{\rm GW}$ in the BDI/CDI/NDI case compared to the AD case. The curvature perturbation $\zeta\propto\omega/k$ in BDI/CDI/NDI around the horizon crossing $k\tau\sim 1$, which is parametrically suppressed compared to the the AD case ($\zeta=O(1)$) for large $k$. This suppression can also be understood from the fact that the matter density is suppressed at early times. Given that GWs are generated mainly around the horizon crossing, $\Omega_{\rm GW}$ in the BDI/CDI/NDI case is suppressed by $(\omega/k)^{4}\propto\zeta^{4}|_{k\tau\sim 1}$. Therefore, it matches the common intuition that isocurvature induced GWs are suppressed because the curvature perturbation in these cases are suppressed.
However, this intuition does not capture all interesting features of DRDI induced GWs. Due to the net anisotropic stress from free-streaming DR, the superhorizon evolution of metric perturbations ($\Phi$ and $\Psi$) is dramatically different from that of $\zeta$ (see figure 1). Therefore, unlike the previous cases, the source term cannot be simply rewritten in terms of $\zeta$. For example, the source term is not suppressed at $\tau\to 0$ even though $\zeta\to 0$ (see figure 2). The induced GWs from DRDI are qualitatively different from other isocurvature cases due to the presence of anisotropic stress, and the density of induced GWs can be much larger than those from CDI since DR density ratio ($R_{\rm dr}$) stay fixed in the radiation domination. Although the usual intuition $\Omega_{\rm GW}\propto\zeta^{4}|_{k\tau\sim 1}\propto R_{\rm dr}^{4}$ (for $R_{\rm dr}\ll 1$) still parametrically holds, it totally obscures the superhorizon dynamics. Therefore, one needs to study the full source term to understand induced GWs for general isocurvature cases.
@@ 4 Constraints from Pulsar Timing Arrays
Having studied the GW spectrum induced by primordial adiabatic and various isocurvature perturbations, we will perform data fitting with the NANOGrav 15-year dataset [NANOGrav:2023gor, NANOGrav:2023hde, NANOGrav:2023hvm, nano_grav_2023_dataset]. In this section, we will present best fit and constraints for AD, CDI and DRDI assuming only one mode dominates for each case. Since BDI and NDI have similar observational signals as CDI, their results can be simply recast from CDI with an overall rescaling. For our data analysis, we employ the parameter $k_{0}$ and the power spectrum amplitude $A_{\rm ad/iso}$ for each mode. We apply the Bayesian inference method to determine the best fit of SIGWs. We adopt PTArcade [andrea_mitridate_2023, Mitridate:2023oar] to sample the posterior probability and generate upper limits above which the additional model is “strongly disfavored” according to the Jeffreys scale [NANOGrav:2023hvm, Jeffreys]. The priors of parameters follow log-uniform distributions within $\log_{10}(k_{0}/{\rm Mpc}^{-1})\in[4,10]$, $\log_{10}(A_{\rm ad})\in[-3,1]$, $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\in[-3,1]$, and $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\in[-3,1]$.
We first show the constraints and best fit for adiabatic (curvature) perturbations (see figure 4). Constraints on curvature perturbations from the latest PTA data have been discussed in several recent studies [NANOGrav:2023hvm, Ellis:2023oxs, Cai:2023dls, Figueroa:2023zhu, Yi:2023tdk, Yi:2023mbm, You:2023rmn, Franciolini:2023pbf, Liu:2023ymk, Liu:2023pau, Liu:2023hpw, Wang:2023ost, Kumar:2024hsi, Chen:2024fir]. Our results with a delta function spectrum agree with those existing results, which serves as the validation of our data analysis. We then present new results for adiabatic perturbations with the broken power law spectrum (see Eq. (2.23)). Current data indicate that the amplitude of the adiabatic power spectrum $A_{\rm ad}\lesssim 10^{-2}$ around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$ for the delta function power spectrum, or $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ for the broken power law spectrum.
The results for CDI and DRDI with the delta function power spectrum exhibit similar features as the AD case. They all have the strongest constraining power around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$, but the sensitivity drops sharply for $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ and gradually decreases for $k_{0}\gtrsim 10^{6}\,\textrm{Mpc}^{-1}$. This reflects the shape of induced GW spectrum (figure 3), which has a peak around $k_{0}$, a sharp edge for large $k$ and a long tail for small $k$. The differences are the peak constraint reads $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\lesssim-2.05$ for CDI and $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\lesssim-1.57$ for DRDI.
@@
The best fit for all cases with delta function power spectrum has a peak at $k_{0}\gg 10^{6}\,\textrm{Mpc}^{-1}$, which indicates that the NANOGrav data only probe the low-$k$ tail of the induced GW spectrum. We show the best fit of the DRDI case as an example in figure 6. As clearly seen in this figure, the peak of the best fit of the delta function power spectrum lies beyond the range of the plot. Interestingly, the posterior of the broken power law case has a bimodal feature. Both AD and DRDI cases have similar features but DRDI is more pronounced. We show the spectrum of GWs from two peaks of the posterior of DRDI in figure 6 with BPL1 ($\log_{10}(k_{0}/\textrm{Mpc}^{-1})=6.82,\log_{10}(R_{\rm dr}^{2}A_{\rm iso})=-1.00$) and BPL2 ($\log_{10}(k_{0}/\textrm{Mpc}^{-1})=7.46,\log_{10}(R_{\rm dr}^{2}A_{\rm iso})=-0.48$). This bimodal feature originates from the fact that the induced GW spectrum for DRDI exhibits “two steps” around the transition scale $k_{0}$ (see figure 3). In particular, the BPL1 with a lower $k_{0}$ fits the data around $f\sim 10^{8}\rm Hz$ well, but not for lower frequencies. In contrast, the BPL2 fits the low frequency data better but misses the features of the data around $f\sim 10^{8}\rm Hz$. It will be interesting to study future high precision data to distinguish these two models. Note that this bimodal feature originates from the sharp transition in the broken power law spectrum at $k_{0}$. A smoother transition will obscure this feature by smearing the induced GW spectrum.
We further compare the constraints on the isocurvature power spectrum from induced GWs with other observations such as the CMB+BAO, Lyman-$\alpha$ forest and CMB spectral distortions. We can see from figure 7 that our constraint from NANOGrav 15 year data places a stringent constraint around $k_{0}\sim 10^{6}\textrm{Mpc}^{-1}$, which are complementary to other constraints on larger scales. We acknowledge that a few other strong constraints on isocurvature are not present in this figure because they are derived assuming either BDI or CDI (see e.g., [Passaglia:2021jla, Graham:2024hah, Inomata:2018htm, Bagherian:2025puf]).
Figure 6: The spectrum of DRDI induced GWs for the best fit parameters with NANOGrav 15 year data. The result for the delta function power spectrum is shown in green. Since the case with the broken power law spectrum has a bimodal feature, we show the results for these two peaks and label them as BPL1 (blue) and BPL2 (orange). The data with error bars from NANOGrav are shown in gray.
@@
Figure 7: Constraints on isocurvature power spectrum for DRDI, CDI and BDI cases with two general parametrizations: a delta function (left, Eq. (2.22)) and a broken power law (right, Eq. (2.23)). Our limit from NANOGrav 15 year data is shown in blue. Gray shaded regions are excluded by the CMB, BAO, Lyman-$\alpha$ forest and CMB spectral distortions [Buckley:2025zgh]. The dotted line indicates the projected sensitivity for the future experiment PIXIE probing CMB distortions [Kogut:2011xw].

### v2

@@ 3 Isocurvature Induced Gravitational Waves
$f(\tau,q,|\mathbf{k}-\mathbf{q}|)$ $=$ $\Phi(q,\tau)\Phi(|\mathbf{k}-\mathbf{q}|,\tau)~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{without DR})$
$+\frac{1}{2}\left(\Phi(q,\tau)+\frac{\partial_{\tau}\Phi(q,\tau)}{\cal H}\right)\left(\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{\partial_{\tau}\Phi(|\mathbf{k}-\mathbf{q}|,\tau)}{\cal H}\right)~.$
Here $c$ and $\Psi,\Phi$ are defined in Eq. (2.9) and initial conditions for $\Psi,\Phi$ are given in section 2. We have used the relation in Eq. (A.2) to get the last line of the above equation.
For the case of DRDI, the source term of Eq. (3.3) becomes
[…]
$\left.+\partial_{a}\Phi\partial_{b}\Psi+\partial_{b}\Phi\partial_{a}\Psi+8\pi Ga^{2}\left(\frac{4}{3}\bar{\rho}_{\gamma_{*}}\partial_{a}v_{\gamma_{*}}\partial_{b}v_{\gamma_{*}}+\bar{\rho}_{\rm dr}\Pi^{(2)}_{\textrm{dr}\,ab}\right)\right$]~, (3.8)
We note that this expression can be reduced to Eq. (3.4) by setting $\Psi=\Phi$ and removing the DR contribution. Since we focus on $R_{\rm dr}\ll 1$ in this study, we only keep the leading terms for $R_{\rm dr}\ll 1$. According to appendix C, the leading contribution of ${\Pi}^{(2)}_{\textrm{dr}\,ij}$ in the Fourier space is given as
${\Pi}^{(2)}_{\textrm{dr}\,ij}(\mathbf{k},\tau)$ $=-\int\frac{\mathrm{d}^{3}\mathbf{q}}{(2\pi)^{3}}{q}_{i}{q}_{j}\int_{0}^{\tau}\mathrm{d}\tilde{\tau}\frac{32}{15}v_{\rm dr}(\mathbf{q},\tilde{\tau})\Phi(\mathbf{k}-\mathbf{q},\tilde{\tau})~~~~~~~~~~(\textrm{DRDI with}~R_{\rm dr}\ll 1)$
$\times\left[j_{0}(k(\tilde{\tau}-\tau))+\frac{10}{7}j_{2}(k(\tilde{\tau}-\tau))+\frac{3}{7}j_{4}(k(\tilde{\tau}-\tau))\right]~,$ (3.9)
where $j_{n}$ is the spherical Bessel function. Here we only keep the terms that are not vanishing after applying the polarization tensor $e^{ij}_{\mathbf{k},\lambda}$.
For the case of DRDI and $R_{\rm dr}\ll 1$, the source term in Eq. (3.6) is now given as
$f(\tau,q,|\mathbf{k}-\mathbf{q}|)=\frac{1}{2}\Phi(q,\tau)\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{1}{2}\Psi(q,\tau)\Psi(|\mathbf{k}-\mathbf{q}|,\tau)$
$~~+\frac{1}{2}\left(\Phi(q,\tau)+\frac{\partial_{\tau}\Psi(q,\tau)}{\cal H}\right)\left(\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{\partial_{\tau}\Psi(|\mathbf{k}-\mathbf{q}|,\tau)}{\cal H}\right)$
$~~+{2R_{\rm dr}^{2}{\cal H}^{2}}v_{\rm dr}(q,\tau)v_{\rm dr}(|\mathbf{k}-\mathbf{q}|,\tau)~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{DRDI})$
$~~+{R_{\rm dr}}{\cal H}\left(\Phi(q,\tau)+\frac{\partial_{\tau}\Psi(q,\tau)}{\cal H}\right)v_{\rm dr}(|\mathbf{k}-\mathbf{q}|,\tau)$
$~~+{R_{\rm dr}}{\cal H}v_{\rm dr}(q,\tau)\left(\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{\partial_{\tau}\Psi(|\mathbf{k}-\mathbf{q}|,\tau)}{\cal H}\right)$
$~~-\frac{16}{5}R_{\rm dr}{\cal H}^{2}\int_{0}^{\tau}d\tilde{\tau}\,v_{\rm dr}(q,\tilde{\tau})\Phi(|\mathbf{k}-\mathbf{q}|,\tilde{\tau})\left[j_{0}(k(\tilde{\tau}-\tau))+\frac{10}{7}j_{2}(k(\tilde{\tau}-\tau))+\frac{3}{7}j_{4}(k(\tilde{\tau}-\tau))\right],$ (3.10)
where the second to fifth lines come from the term $\propto\partial_{i}v_{\gamma*}\partial_{j}v_{\gamma*}$, while the last term is from $\Pi_{\textrm{dr},ij}^{(2)}$. Since the transfer functions $\Psi,\Phi$ is $\mathcal{O}(R_{\rm dr})$ and $v_{\rm dr}$ is $\mathcal{O}(R_{\rm dr}^{0})$ in DRDI, all terms in the above equation are of the same order $\mathcal{O}(R^{2}_{\rm dr})$.
The energy density of gravitational waves relevant for observations is given by [32]
@@
where $f(x,u,v)$ is given in Eqs. (3) or (3.10). Plugging in the Green’s function of gravitational waves, the kernel function can be further split as
@@
where $h\equiv H_{0}/(100\,\textrm{km/s/Mpc})$, $g_{*}(T)$ and $g_{*S}(T)$ are the effective number of relativistic degrees of freedom for energy and entropy, respectively. $T_{0}$ is the present-day temperature and $T_{\rm GW}$ is the temperature when the GWs are generated, which is $T_{\rm GW}=O(100)$MeV in this study. Therefore, $g_{*}(T_{\rm GW})=g_{*S}(T_{\rm GW})=17.25$, which includes the contribution from pions, muons, and electrons/positions. The radiation density today $\Omega_{r,0}h^{2}=4.2\times 10^{-5}$ is taken from [1].
To explicitly view the effects on induced GWs from different initial conditions, we should look at the source term, in particular $f(x,u,v)$ defined in Eqs. (3) or (3.10) which encodes all the time evolution. As can be seen from figure 2, the subhorizon dynamics are rather universal, which exhibit damped oscillations. However, the superhorizon behavior depends on the certain initial condition. Both the adiabatic and DRDI cases remain constant at superhorizon $x\ll 1$, whereas the CDI case exhibits a $\tau^{2}$ dependence. These features originate from the fact that $\Phi$ and $\Psi$ are constant for AD and DRDI but $\propto\tau$ for CDI outside the horizon. We note that the CDI case is further suppressed by $(k_{\rm eq}/k)^{2}$ and the DRDI case by $R_{\rm dr}^{2}$, which reflects the suppression of the density of the corresponding species. Therefore, for modes relevant for PTA observations ($k\gg k_{\rm eq}$), the source term of CDI is highly suppressed but that of DRDI can still be sizable for appropriate $R_{\rm dr}^{2}$. The cases of BDI and NDI are similar to CDI with an overall rescaling, and thus we don’t show them here.
@@
Here we would like to clarify some intuitions for curvature and isocurvature induced GWs. One common intuition of SIGWs is that GWs are always sourced by metric perturbations (or curvature perturbations). This is indeed the case for previous studies where there is no anisotropic stress ($\Phi=\Psi$) and the source term of GWs is simply a function of $\Phi$ and its derivatives (see Eq. (3)). The size of curvature and isocurvature induced GWs can be estimated by the peak value of $\Phi$, which is around the horizon crossing: $\Omega_{\rm GW}\propto\Phi^{4}|_{k\tau\sim 1}$. Therefore, it is intuitive to understand that GWs induced by isocurvature cases without anisotropic stress (CDI/BDI/NDI) are suppressed (compared to the AD case) because $\Phi^{4}|_{k\tau\sim 1}\propto(\omega/k)^{4}$ for $k\gg\omega$. This suppression can also be understood from the fact that the matter density is suppressed at early times.
However, this intuition does not capture all interesting features of DRDI induced GWs. Due to the net anisotropic stress from free-streaming DR, the source term cannot be simply rewritten in terms of metric perturbations $\Phi$ and $\Psi$. There are additional terms involving $R_{\rm dr}v_{\rm dr}$ (see Eq. (3.10)). These terms contribute the same order as metric perturbations because $\Phi,\Psi$ is $\mathcal{O}(R_{\rm dr})$ in DRDI with $R_{\rm dr}\ll 1$. Therefore, the spectrum of DRDI induced GWs is qualitatively different from other isocurvature cases, and the density of induced GWs can be much larger than those from CDI since DR density ratio ($R_{\rm dr}$) stay fixed in the radiation domination. Although the usual intuition $\Omega_{\rm GW}\propto\Phi^{4}|_{k\tau\sim 1}\propto R_{\rm dr}^{4}$ (for $R_{\rm dr}\ll 1$) still parametrically holds, it totally obscures the novel contribution from DR density perturbations. Therefore, one needs to study the full source term to understand induced GWs for general isocurvature cases.
@@ 4 Constraints from Pulsar Timing Arrays
Having studied the GW spectrum induced by primordial adiabatic and various isocurvature perturbations, we will perform data fitting with the NANOGrav 15-year dataset [28, 85, 86, 87]. In this section, we will present best fit and constraints for AD, CDI and DRDI assuming only one mode dominates for each case. Since BDI and NDI have similar observational signals as CDI, their results can be simply recast from CDI with an overall rescaling. For our data analysis, we employ the parameter $k_{0}$ and the power spectrum amplitude $A_{\rm ad/iso}$ for each mode. We apply the Bayesian inference method to determine the best fit of SIGWs. We adopt PTArcade [88, 89] to sample the posterior probability and generate upper limits above which the additional model is “strongly disfavored” according to the Jeffreys scale [86, 90]. The priors of parameters follow log-uniform distributions within $\log_{10}(k_{0}/{\rm Mpc}^{-1})\in[4,10]$, $\log_{10}(A_{\rm ad})\in[-3,1]$, $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\in[-3,1]$, and $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\in[-3,1]$.
We first show the constraints and best fit for adiabatic (curvature) perturbations (see figure 4). Constraints on curvature perturbations from the latest PTA data have been discussed in several recent studies [86, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103]. Our results with a delta function spectrum agree with those existing results, which serves as the validation of our data analysis. We then present new results for adiabatic perturbations with the broken power law spectrum (see Eq. (2.23)). Current data indicate that the amplitude of the adiabatic power spectrum $A_{\rm ad}\lesssim 10^{-2}$ around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$ for the delta function power spectrum, or $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ for the broken power law spectrum.
The constraint for CDI and DRDI with the delta function power spectrum exhibit similar features as the AD case. They all have the strongest constraining power around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$, but the sensitivity drops sharply for $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ and gradually decreases for $k_{0}\gtrsim 10^{6}\,\textrm{Mpc}^{-1}$. This reflects the shape of induced GW spectrum (figure 3), which has a peak around $k_{0}$, a sharp edge for large $k$ and a long tail for small $k$. The differences are the peak constraint reads $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\lesssim-2.05$ for CDI and $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\lesssim-1.38$ for DRDI.
@@
The $1\sigma$ ($2\sigma$) contours shown in figures 4 and 5 reveal a correlation between the amplitude of the power spectrum and the characteristic scale $k_{0}$. This behavior reflects the degeneracy between these two parameters when data are mostly sensitive to the IR portion of the induced GW spectrum. We also show the spectrum of induced GWs (in figure 6) for best fit parameters of different cases listed in table 1. As seen in this figure, the peaks of the spectrum of AD and CDI cases lie beyond the range of data, meaning the data only probe the IR part of the spectrum. Interestingly, the peak of the DRDI induced GWs sits within the data because the spectrum near the peak captures some interesting features of the data around $f\sim 10^{8}$ Hz. This difference originates from the peculiar shape of DRDI induced GWs shown in figure 3. For both choices of the isocurvature power spectrum, the data exhibit mild preference for DRDI induced GWs over AD and CDI cases. It will be interesting to study future high precision data to distinguish different isocurvature scenarios.
Delta function Broken power law
AD DRDI CDI AD DRDI CDI
$\log_{10}(k_{0}/{\rm Mpc}^{-1})$ 9.69 6.97 7.69 7.39 6.94 8.12
Rescaled Amplitude -0.0523 -1.08 -1.27 -1.30 -0.879 -0.668
Table 1: Best fit parameters of the amplitude of the power spectrum and characteristic scale $k_{0}$ for AD, CDI and DRDI cases. The value of the rescaled amplitude is $\log_{10}(A_{\rm ad})$ for AD, $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})$ for DRDI and $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})$ for CDI.
We further compare the constraints on the isocurvature power spectrum from induced GWs with other observations such as the CMB+BAO, Lyman-$\alpha$ forest and CMB spectral distortions. We can see from figure 7 that our constraint from NANOGrav 15 year data places a stringent constraint around $k_{0}\sim 10^{6}\textrm{Mpc}^{-1}$, which are complementary to other constraints on larger scales. We acknowledge that a few other strong constraints on isocurvature are not present in this figure because they are derived assuming either BDI or CDI (see e.g., [104, 63, 21, 26]).
Figure 6: The spectrum of induced GWs for the best fit parameters with NANOGrav 15 year data. We show the results for two choices of isocurvature power spectrum: a delta function (top, Eq. (2.22)) and a broken power law (bottom, Eq. (2.23)). The data with error bars from NANOGrav are shown in gray.
@@
Figure 7: Constraints on isocurvature power spectrum for DRDI, CDI and BDI cases with two general parametrizations: a delta function (left, Eq. (2.22)) and a broken power law (right, Eq. (2.23)). Our limit from NANOGrav 15 year data is shown in blue. Gray shaded regions are excluded by the CMB, BAO, Lyman-$\alpha$ forest and CMB spectral distortions [22]. The dotted line indicates the projected sensitivity for the future experiment PIXIE probing CMB distortions [105].